OpenStudy (anonymous):

Please help?? A sample of a gas has a volume of 1.5 L at 150 K. If the gas is heated to 235 K at constant pressure, what will its final volume be?

OpenStudy (aaronq):

use the ideal gas law, PV=nRT

OpenStudy (julian.pitalua):

Because the pressure is constant , we have got that temperature and volume are directly related. If the temperature rises , then the volume will too. T1/V1=T2/V2 150K/1,5L= 235K/ xx L

OpenStudy (julian.pitalua):

Also we assume that the number of moles is constant